OpenStudy (samsan9):

For f(x)=2x+1 and g(x)=x^2-7, find (f-g)(x)

OpenStudy (anonymous):

any thoughts already?

OpenStudy (samsan9):

I thinkthat It might be -x^2+2x-6 and it is one of the answer choices

OpenStudy (samsan9):

the answer choices are a) x^2-2x-8 b.)2x^2-15 c.)-x2+2x-6 d.)8+2x-12x

OpenStudy (anonymous):

OK but why -6? isn't it 8?

OpenStudy (samsan9):

sorry my laptop is buggy and so are you saying that it is a?

OpenStudy (jdoe0001):

\(\bf f(x)=2x+1 \qquad g(x)=x^2-7 \\ \quad \\ (f-g)(x)\to f(x)-g(x)\to (2x+1)-(x^2-7)\)

OpenStudy (samsan9):

ok so I did that and I got the letter choice c unless I did something wrong

OpenStudy (anonymous):

Is \(g=x^2+7\)?

OpenStudy (jdoe0001):

\(\bf f(x)=2x+1 \qquad g(x)=x^2-7 \\ \quad \\ (f-g)(x)\to f(x)-g(x)\to (2x+1)-(x^2-7) \\ \quad \\ \to 2x+1{\color{brown}{ -x^2+7}}\to -x^2+2x-8\)

OpenStudy (jdoe0001):

hhmm \(\bf f(x)=2x+1 \qquad g(x)=x^2-7 \\ \quad \\ (f-g)(x)\to f(x)-g(x)\to (2x+1)-(x^2-7) \\ \quad \\ \to 2x+1{\color{brown}{ -x^2+7}}\to -x^2+2x+8\) rather

OpenStudy (jdoe0001):

anyhow.. the -() in front of g(x) turns all signs the other way, thus

OpenStudy (samsan9):

so the answer is D

OpenStudy (jdoe0001):

hmm I'd think you have a typo on D, so... I guess so =)

OpenStudy (samsan9):

thank you :D

OpenStudy (jdoe0001):

yw